Mabel's salary starts at $16,000 per year with annual raises of $1500. Janet's starting salary is $19,300 with annual raises of $950. After how many years will the two women be earning the same salary?
Answer by Maths68(1474)

Let
It takes = x years

Mabel's slary 16,000
Annual raise = 1500

Janet's Salary = 19300
Annual raise = 950

16000+ 1500*x = 19300 + 950*x
1500x-950x=19300-16000
550x=3300
550x/550=3300/550
x=6


In 6 years they will be earning same salary.

Check
======
16000+ 1500*x = 19300 + 950*x
16000 + 1500 * 6 = 19300 + 950*6
16000 + 9000 = 19300+5700
25000 = 25000